OpenStudy (anonymous):

Simplify this expression. 9x(5x^2)

OpenStudy (diyadiya):

What is\( ~~~9 \times 5 \)??

OpenStudy (anonymous):

45

OpenStudy (diyadiya):

ok so \( 9x(5x^2) = 45 x.x^2\) what is \( x \times x^2 \) ??

OpenStudy (anonymous):

\[2x^2\]?

OpenStudy (diyadiya):

No! \( \Large x^m \times x^n = x^{m+n} \) \( \Large x^1 \times x^2 = ?? \) remember: x is same as \( x^1\)

OpenStudy (anonymous):

so it's \[x^3\]

OpenStudy (diyadiya):

Yes so final answer is ?

OpenStudy (anonymous):

\[45x^3\]

OpenStudy (diyadiya):

Right! :)

OpenStudy (anonymous):

Thanks for your help I actually understand now!

OpenStudy (diyadiya):

No problem :D